# Ceresa cycles of bielliptic Picard curves

### v1

@@ 5.5. Proof of the main theorem
Putting everything together, we can prove the following theorem which implies Theorem 1.1 of the introduction.
@@
Let $k$ be a field of characteristic $\neq 2,3$ and let $C/k$ be a smooth projective curve with equation $y^{3}=x^{4}+ax^{2}+b$ and Jacobian $J$. Then the Ceresa cycle $\kappa(C)\in{\rm CH}^{2}(J)$ with respect to the point at infinity is torsion if and only if $\mathcal{Q}_{(a,b)}=((a^{2}-4b),a(a^{2}-4b))$ is a torsion point of the elliptic curve $E^{\Delta}\colon y^{2}=x^{3}+4b(a^{2}-4b)^{2}$.
@@
The curve $C$ is isomorphic to the pullback $\mathcal{C}_{s}$ of $\mathcal{C}\rightarrow S$ along the $k$-point $s=(a,b)\colon{\rm Spec}(k)\rightarrow S$. By Corollary 5.13, $\kappa(C)$ is torsion if and only if $\sigma(s)\in\mathcal{E}^{\Delta}_{s}(k)$ is torsion. By Theorem 5.15, $\sigma(s)$ is a multiple of $\mathcal{Q}_{s}=\mathcal{Q}_{(a,b)}$, so $\sigma(s)$ is torsion if and only if $\mathcal{Q}_{s}=(a^{2}-4b,a(a^{2}-4b))$ is torsion. ∎
@@
We take $(a,b)=(2t,1)$. A simple computation (using that $k=\overline{k}$) shows that there is an isomorphism from $E_{a,b}^{\Delta}$ to $\widehat{E}\colon y^{2}=x^{3}+1$, sending $\mathcal{Q}_{(a,b)}=(4t^{2}-4,2t(4t^{2}-4))$ to $Q_{t}=(\sqrt[3]{t^{2}-1},t)$, for some choice of cube root. Thus $\kappa(C_{t})$ is torsion if and only if $Q_{t}$ is torsion. ∎
@@
Suppose $C$ is a nice curve over ${\rm\mathbb{Q}}$ that is ${\rm\overline{\mathbb{Q}}}$-isomorphic to some bielliptic Picard curve $C_{a,b}$ and has torsion Ceresa cycle. Then $C$ is ${\rm\overline{\mathbb{Q}}}$-isomorphic to either $C_{0}$, $C_{3}$, or $C_{\sqrt{-3}}$.
@@
Recall from §5.3 that $\Phi_{S}(\kappa(\mathcal{C})_{(1)})=\sigma/d$ in $\mathcal{E}^{\Delta}(S)\otimes{\rm\mathbb{Q}}$, where $d$ is the denominator of $\sigma_{0}$. Moreover $\sigma=N\cdot\mathcal{Q}$ in $\mathcal{E}^{\Delta}(S)$ by Theorem 5.15. So $d\cdot\Phi_{S}(\kappa(\mathcal{C})_{(1)})-N\cdot\mathcal{Q}$ is torsion in $\mathcal{E}^{\Delta}(S)$, so
(5.6) $M\cdot(d\cdot\Phi_{S}(\kappa(\mathcal{C})_{(1)})-N\cdot\mathcal{Q})=0\text{ in }\mathcal{E}^{\Delta}(S)$
for some $M\geq 1$. The orders of the torsion specializations $\mathcal{Q}_{s}\in\mathcal{E}^{\Delta}_{s}({\rm\mathbb{C}})$, where $s$ varies in $S({\rm\mathbb{C}})$, are unbounded. The relation (5.6) implies that the same is true for the orders of the torsion specializations of $\kappa(\mathcal{C}_{s})_{(1)}$, which equals $\kappa(\mathcal{C}_{s})$ by Lemma 5.12. ∎
@@
$\langle\,,\,\rangle_{\mathrm{BB}}\colon{\rm CH}^{i}(A)_{\mathrm{hom}}\times{\rm CH}^{g+1-i}(A)_{\mathrm{hom}}\to{\rm\mathbb{R}}$
has been constructed unconditionally by Künnemann [27, Corollary 1.7].
@@
where $\langle\,,\,\rangle_{\mathrm{NT}}$ is the Neron–Tate height pairing on $E^{\Delta}({\rm\overline{\mathbb{Q}}})$.
@@
This follows from Theorem 5.15 and Lemma 5.12, using the fact that the Beilinson–Bloch height is compatible with correspondences [6, 4.0.3] and agrees with the Neron–Tate height pairing for divisors on curves [6, 4.0.8]. ∎
@@
The theorem follows from Theorem 5.19 and the nondegeneracy and Northcott properties of the Neron–Tate height of an elliptic curve. ∎
@@ 6. Algebraic triviality and the Griffiths group
Let $C$ be a nice curve over a field $k$ and $J$ be its Jacobian. The Griffiths group of $1$-cycles on $J$ is the group ${\rm Gr}_{1}(J)={\rm Gr}^{g-1}(J)={\rm CH}_{1}(J)_{0}/A_{1}(J)$, where $A_{1}(J)$ is the subgroup of algebraically trivial $1$-cycles. The very general Ceresa cycle (of a genus $g\geq 3$ curve over ${\rm\mathbb{C}}$) is known to be of infinite order in ${\rm Gr}_{1}(J)$ [35]. In particular, the Ceresa cycle of a very general curve is not algebraically trivial, nor is any multiple of it. In this section we show that this fails for bielliptic Picard curves. We also discuss the Beilinson–Bloch conjectures in this context, and show that they imply that ${\rm Gr}^{2}(J)$ is finite for certain bielliptic Picard curves.
@@
There exists an integer $N$, a smooth projective relative curve $\mathcal{X}\rightarrow S$ with two sections $\sigma_{1},\sigma_{2}$ and a cycle $\mathcal{Z}$ on ${\rm CH}^{1}(\mathcal{J}\times_{S}X)$, with the property that $\mathcal{Z}_{\sigma_{1}}=N\cdot\kappa(\mathcal{C})_{(1)}$ and $\mathcal{Z}_{\sigma_{2}}=0$.
This proposition implies Theorem 1.5, since for every $k$-point $s\colon{\rm Spec}(k)\rightarrow S$, corresponding to the bielliptic Picard curve $\mathcal{C}_{s}=C$ with $X=\mathcal{X}_{s}$, $p_{i}=\sigma_{i}(s)$ and $Z=\mathcal{Z}_{s}$, the cycle $Z\in{\rm CH}^{1}(J\times_{k}X)$ has the property that $Z_{p_{1}}=N\cdot\kappa(C)_{(1)}$ and $Z_{p_{2}}=0$. Since $\kappa(C)_{(1)}=\kappa(C)$ by Lemma 5.12, $N\cdot\kappa(C)$ is algebraically equivalent to zero.
@@
We will use the isomorphisms $\Phi$ from Corollary 5.11. Recall from §5.3 that there exists an $N_{1}$ such that $\Phi_{S}(N_{1}\cdot\kappa(\mathcal{C})_{(1)})=\sigma\in\mathcal{E}^{\Delta}(S)$. We can deform the section $\sigma$ of $\mathcal{E}^{\Delta}$ to the identity section $\infty\colon S\rightarrow\mathcal{E}^{\Delta}$. More precisely, there exists a smooth projective curve $\mathcal{X}\rightarrow S$, a section $\mathcal{V}\in\mathcal{E}^{\Delta}(\mathcal{X})$ and sections $\sigma_{1},\sigma_{2}$ of $\mathcal{X}\rightarrow S$, such that $\mathcal{V}_{\sigma_{1}}=\sigma$ and $\mathcal{V}_{\sigma_{2}}=\infty$. In fact, $\mathcal{X}=\mathcal{E}^{\Delta}$ with $\sigma_{1}=\sigma$, $\sigma_{2}=\infty$ and with $\mathcal{V}=\Delta_{\mathcal{X}/S}\subset\mathcal{X}\times_{S}\mathcal{X}=\mathcal{E}^{\Delta}\times_{S}\mathcal{X}$ will do. There exists an $N_{2}$ such that $\Phi_{\mathcal{X}}^{-1}(N_{2}\cdot\mathcal{V})\in{\rm CH}^{2}_{(1)}(\mathcal{J}\times_{S}\mathcal{X};{\rm\mathbb{Q}})$ is the image of an integral cycle $\mathcal{W}\in{\rm CH}^{2}(\mathcal{J}\times_{S}\mathcal{X})$. The commutativity of the diagram of Corollary 5.11 applied to $\mathcal{X}\rightarrow S$ (which is flat, so admits pullbacks) shows that $\mathcal{W}_{\sigma_{1}}=N_{1}N_{2}\cdot\kappa(\mathcal{C})$ and $\mathcal{W}_{\sigma_{2}}=0$ in ${\rm CH}^{2}(\mathcal{J};{\rm\mathbb{Q}})$. In other words, the elements $\mathcal{W}_{\sigma_{1}}-N_{1}N_{2}\kappa(\mathcal{C})$ and $\mathcal{W}_{\sigma_{2}}$ are torsion in ${\rm CH}^{2}(\mathcal{J})$. There exists an integer $N_{3}$ such that these elements are killed by $N_{3}$, so the elements $N=N_{1}N_{2}N_{3}$ and $\mathcal{Z}=N_{3}\cdot\mathcal{W}$ satisfy the conclusion of the proposition. ∎
@@ 6.2. Beilinson–Bloch conjectures
We briefly describe some consequences of the Beilinson–Bloch conjecture [6, 8] in our setting. Let $X$ be a nice variety over a number field $k$. Fix a prime $\ell$ and denote the $\ell$-adic cohomology $\operatorname{H}^{i}(X_{\bar{k}},{\rm\mathbb{Q}}_{\ell})$ simply by $\operatorname{H}^{i}(X)$. The conjecture predicts that for each $i\in{\rm\mathbb{Z}}$, the group of homologically trivial cycles ${\rm CH}^{i}(X)_{\mathrm{hom}}$ is finitely generated, the $L$-function $L(\mathfrak{h}^{2i-1}(X),s)$ has analytic continuation and functional equation, and
@@ 6.3. The Griffiths group
Using the isogeny decomposition $J\sim P\times E$ and the Kunneth formula, one checks that the first graded piece $\mathrm{gr}^{0}\operatorname{H}^{3}(J)(2)$ in the coniveau filtration is a nonzero quotient of $V\otimes\operatorname{H}^{1}(E)(2)$. Define
@@
The curve $C_{0,1}\colon y^{3}=x^{4}+1$ has CM by ${\rm\mathbb{Q}}(i)$ [28, Example 6.15]. In [30, Theorem 1.3], Lilienfeldt and the second author show that the corresponding Hecke $L$-function has nonvanishing central value, and hence ${\rm Gr}^{2}(J_{0,1})$ is conjecturally finite.

### v2

@@ 5.5. Proof of the main theorem
Putting everything together, we prove the following theorem which implies Theorem 1.1 of the introduction.
@@
Let $k$ be a field of characteristic $\neq 2,3$ and let $C/k$ be a smooth projective curve with equation $y^{3}=x^{4}+ax^{2}+b$ and Jacobian $J$. Then the Ceresa cycle $\kappa_{\infty}(C)\in{\rm CH}^{2}(J)$ is torsion if and only if $\mathcal{Q}_{(a,b)}=((a^{2}-4b),a(a^{2}-4b))$ is a torsion point of the elliptic curve $E^{\Delta}\colon y^{2}=x^{3}+4b(a^{2}-4b)^{2}$.
Moreover, the quotient of the torsion orders $\operatorname*{ord}(\kappa_{\infty}(C))/\operatorname*{ord}(\mathcal{Q}_{a,b})$, whenever it is defined, takes finitely many values when varying over all $k$ and $C$ as above.
@@
The curve $C$ is isomorphic to the pullback $\mathcal{C}_{s}$ of $\mathcal{C}\rightarrow S$ along the $k$-point $s=(a,b)\colon{\rm Spec}(k)\rightarrow S$. By Corollary 5.13, $\kappa_{\infty}(C)$ is torsion if and only if $\sigma(s)\in\mathcal{E}^{\Delta}_{s}(k)$ is torsion. By Theorem 5.16, $\sigma(s)$ is a multiple of $\mathcal{Q}_{s}=\mathcal{Q}_{(a,b)}$, so $\sigma(s)$ is torsion if and only if $\mathcal{Q}_{s}=(a^{2}-4b,a(a^{2}-4b))$ is torsion. Tracing through these equivalences and using Lemma 5.14, the quotients of the torsion orders take finitely many values at each step, so $\operatorname*{ord}(\kappa_{\infty}(C))/\operatorname*{ord}(\mathcal{Q}_{a,b})$ takes finitely many values as well. ∎
@@
We take $(a,b)=(2t,1)$. A simple computation (using that $k=\overline{k}$) shows that there is an isomorphism from $E_{a,b}^{\Delta}$ to $\widehat{E}\colon y^{2}=x^{3}+1$, sending $\mathcal{Q}_{(a,b)}=(4t^{2}-4,2t(4t^{2}-4))$ to $Q=(\sqrt[3]{t^{2}-1},t)$, for some choice of cube root. Thus $\kappa_{\infty}(C_{t})$ is torsion if and only if $Q$ is torsion. ∎
@@
Suppose $C$ is a curve over ${\rm\mathbb{Q}}$ that is ${\rm\overline{\mathbb{Q}}}$-isomorphic to some bielliptic Picard curve $C_{a,b}$ and has torsion Ceresa cycle. Then $C$ is ${\rm\overline{\mathbb{Q}}}$-isomorphic to either $C_{0}$, $C_{3}$, or $C_{\sqrt{-3}}$.
@@
The torsion specializations $\mathcal{Q}_{s}\in\mathcal{E}^{\Delta}_{s}({\rm\mathbb{C}})$, where $s$ varies in $S({\rm\mathbb{C}})$, are all defined over ${\rm\overline{\mathbb{Q}}}$ and have unbounded order. We conclude by Theorem 5.17. ∎
@@
$\langle\,,\,\rangle_{\mathrm{BB}}\colon{\rm CH}^{i}(A;{\rm\mathbb{Q}})_{\mathrm{hom}}\times{\rm CH}^{g+1-i}(A;{\rm\mathbb{Q}})_{\mathrm{hom}}\to{\rm\mathbb{R}}$
has been constructed unconditionally by Künnemann [31, Corollary 1.7].
@@
where $\langle\,,\,\rangle_{\mathrm{NT}}$ is the Néron–Tate height pairing on $E^{\Delta}({\rm\overline{\mathbb{Q}}})$.
@@
This follows from Theorem 5.16 and Lemma 5.12, using the fact that the Beilinson–Bloch height is compatible with correspondences [6, 4.0.3] and agrees with the Néron–Tate height pairing for divisors on curves [6, 4.0.8]. ∎
@@
The theorem follows from Theorem 5.20 and the nondegeneracy and Northcott properties of the Néron–Tate height of an elliptic curve. ∎
@@ 6. Algebraic triviality and the Griffiths group
Let $C$ be a curve over a field $k$ and $J$ be its Jacobian. The Griffiths group of $1$-cycles on $J$ is the group ${\rm Gr}_{1}(J)={\rm Gr}^{g-1}(J)={\rm CH}_{1}(J)_{0}/A_{1}(J)$, where $A_{1}(J)$ is the subgroup of algebraically trivial $1$-cycles. The very general Ceresa cycle (of a genus $g\geq 3$ curve over ${\rm\mathbb{C}}$) is known to be of infinite order in ${\rm Gr}_{1}(J)$ [40]. In particular, the Ceresa cycle of a very general curve is not algebraically trivial, nor is any multiple of it. In this section we show that this fails for bielliptic Picard curves. We also discuss the Beilinson–Bloch conjectures in this context, and show that they imply that ${\rm Gr}^{2}(J)$ is finite for certain bielliptic Picard curves.
@@
Let $\kappa_{{\rm\mathbb{Z}}}\in{\rm CH}^{2}(\mathcal{J})$ be a class and $N_{1}\geq 1$ an integer satisfying the conclusions of Lemma 5.14. Then there exists an integer $N\geq 1$, a smooth projective relative curve $\mathcal{X}\rightarrow S$ with two sections $\sigma_{1},\sigma_{2}$ and a cycle $\mathcal{Z}$ on ${\rm CH}^{1}(\mathcal{J}\times_{S}X)$, with the property that $\mathcal{Z}_{\sigma_{1}}=N\cdot\kappa_{{\rm\mathbb{Z}}}$ and $\mathcal{Z}_{\sigma_{2}}=0$.
This proposition implies Theorem 1.5, since for every $k$-point $s\colon{\rm Spec}(k)\rightarrow S$, corresponding to the bielliptic Picard curve $\mathcal{C}_{s}=C$ with $X=\mathcal{X}_{s}$, $p_{i}=\sigma_{i}(s)$ and $Z=\mathcal{Z}_{s}$, the cycle $Z\in{\rm CH}^{1}(J\times_{k}X)$ has the property that $Z_{p_{1}}=s^{*}(\kappa_{{\rm\mathbb{Z}}})=NN_{1}\cdot\kappa_{\infty}(C)$ and $Z_{p_{2}}=0$. By definition, this means that $(NN_{1})\cdot\kappa_{\infty}(C)$ is algebraically trivial.
@@
We will use the isomorphisms $\Phi$ from Corollary 5.11. Recall from §5.3 that there exists an integer $N_{2}\geq 1$ such that $\Phi_{S}(N_{2}\cdot\kappa(\mathcal{C})_{(1)})=\sigma\in\mathcal{E}^{\Delta}(S)$. We can deform the section $\sigma$ of $\mathcal{E}^{\Delta}$ to the identity section $\infty\colon S\rightarrow\mathcal{E}^{\Delta}$. More precisely, there exists a smooth projective relative curve $\mathcal{X}\rightarrow S$, a section $\mathcal{V}\in\mathcal{E}^{\Delta}(\mathcal{X})$ and sections $\sigma_{1},\sigma_{2}$ of $\mathcal{X}\rightarrow S$, such that $\mathcal{V}_{\sigma_{1}}=\sigma$ and $\mathcal{V}_{\sigma_{2}}=\infty$. In fact, $\mathcal{X}=\mathcal{E}^{\Delta}$ with $\sigma_{1}=\sigma$, $\sigma_{2}=\infty$ and with $\mathcal{V}=\Delta_{\mathcal{X}/S}\subset\mathcal{X}\times_{S}\mathcal{X}=\mathcal{E}^{\Delta}\times_{S}\mathcal{X}$ will do. There exists an $N_{3}\geq 1$ such that $\Phi_{\mathcal{X}}^{-1}(N_{3}\cdot\mathcal{V})\in{\rm CH}^{2}_{(1)}(\mathcal{J}\times_{S}\mathcal{X};{\rm\mathbb{Q}})$ is the image of an integral cycle $\mathcal{W}_{1}\in{\rm CH}^{2}(\mathcal{J}\times_{S}\mathcal{X})$. Let $\mathcal{W}=N_{1}\cdot\mathcal{W}_{1}$. The commutativity of the diagram of Corollary 5.11 applied to $\mathcal{X}\rightarrow S$ (which is flat, so admits pullbacks) shows that $\mathcal{W}_{\sigma_{1}}=N_{1}N_{2}N_{3}\cdot\kappa(\mathcal{C})_{(1)}=N_{2}N_{3}\kappa_{{\rm\mathbb{Z}}}$ and $\mathcal{W}_{\sigma_{2}}=0$ in ${\rm CH}^{2}(\mathcal{J};{\rm\mathbb{Q}})$. In other words, the elements $\mathcal{W}_{\sigma_{1}}-N_{2}N_{3}\kappa_{{\rm\mathbb{Z}}}$ and $\mathcal{W}_{\sigma_{2}}$ are torsion in ${\rm CH}^{2}(\mathcal{J})$. There exists an integer $N_{4}\geq 1$ such that these elements are killed by $N_{4}$, so the elements $N=N_{2}N_{3}N_{4}$ and $\mathcal{Z}=N_{4}\cdot\mathcal{W}$ satisfy the conclusion of the proposition. ∎
@@ 6.2. Beilinson–Bloch conjectures
We briefly describe some consequences of the Beilinson–Bloch conjecture [6, 8] in our setting. Let $X$ be a smooth projective geometrically integral variety over a number field $k$. Fix a prime $\ell$ and denote the $\ell$-adic cohomology $\operatorname{H}^{i}(X_{\bar{k}},{\rm\mathbb{Q}}_{\ell})$ simply by $\operatorname{H}^{i}(X)$. The conjecture predicts that for each $i\in{\rm\mathbb{Z}}$, the group of homologically trivial cycles ${\rm CH}^{i}(X)_{\mathrm{hom}}$ is finitely generated, the $L$-function $L(\mathfrak{h}^{2i-1}(X),s)$ has analytic continuation and functional equation, and
@@ 6.3. The Griffiths group
Using the isogeny decomposition $J\sim P\times E$ and the Künneth formula, one checks that the first graded piece $\mathrm{gr}^{0}\operatorname{H}^{3}(J)(2)$ in the coniveau filtration is a nonzero quotient of $V\otimes\operatorname{H}^{1}(E)(2)$. Define
@@
The curve $C_{0,1}\colon y^{3}=x^{4}+1$ has CM by ${\rm\mathbb{Q}}(i)$ [32, Example 6.15]. In [35, Theorem 1.3], Lilienfeldt and the second author show that the corresponding Hecke $L$-function has nonvanishing central value, and hence ${\rm Gr}^{2}(J_{0,1})$ is conjecturally finite.